OpenStudy (anonymous):

Please help me with an Algebra II question. I will give a medal and become a fan. (Question and answer choices down below...)

OpenStudy (anonymous):

Which radical expression is equivalent to \[5\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

A. \[\sqrt{5}\] B. \[\sqrt{125}\] C. \[\sqrt[3]{5}\] D. \[\sqrt[3]{25}\]

OpenStudy (anonymous):

I believe it is B, but I'm not a hundred percent sure. Sorry

OpenStudy (anonymous):

That's correct! The answer is B. \[\sqrt{125}\]

OpenStudy (anonymous):

@Maddiem

OpenStudy (anonymous):

awesome!